OpenStudy (anonymous):

What is 12x-96 can somebody break it down

OpenStudy (anonymous):

What do you mean by break it down? Do you have a value for x, or does your teacher want you do expand the entire problem?

OpenStudy (anonymous):

Factors the expression by using the pattern for the sum or different of cubes. The problem is 12x to the 3rd power subtract 96

OpenStudy (anonymous):

If we factor, we get \[12(x-8)\]

OpenStudy (phi):

do you mean \[ 12x^3 -96 \] ? if you should write it as 12x^3 -96

OpenStudy (anonymous):

Oh, I didn't notice that exponent. I apologize. Let me resolve

OpenStudy (anonymous):

Factoring \[12x^3-96\] gives us \[12(x-2)(x^2+2x+4)\]

OpenStudy (anonymous):

Yes phi thanks guys!!! Yes that is what I ment

OpenStudy (anonymous):

Yes it does but he wants me to show my work

OpenStudy (phi):

they want you to learn how to factor 12x^3 - 96 the first step is factor 12 out of each term to get \[ 12 (x^3 - 8)\] the next "trick" is to know that 8 = 2*2*2= 2^3 \[ 12 (x^3 - 2^3)\] now use the "pattern" they want you to learn to factor the difference of cubes

OpenStudy (anonymous):

X^3 + 125 @mossyfish @phi

OpenStudy (anonymous):

\[(x+5)(x2−5x+25) \]